OpenStudy (anonymous):

I need help!! (k-1)(k+5)=0

OpenStudy (anonymous):

I need to use the zero product property to solve it, but I don't remember how to do it.

OpenStudy (skullpatrol):

Recall the zero product property: ab = 0 if and only if a=0 or b=0.

OpenStudy (skullpatrol):

Let a=k-1 and b=k+5

OpenStudy (anonymous):

Do I multiply the numbers?

OpenStudy (skullpatrol):

Nope, just solve the two equations: k-1=0 and k+5=0.

OpenStudy (skullpatrol):

Your answer will tell you for which values of k the equation (k-1)(k+5)=0 will be a true statement, Because 0(any number)=0 and vice versa.

OpenStudy (anonymous):

Ok! Thank you so much!

OpenStudy (skullpatrol):

Thanks for trying to learn :-)